OpenStudy (anonymous):

The difference of two numbers, a and b, is 21. The difference of 5 times a and two times b is 18. What are the values of a and b?

OpenStudy (anonymous):

This one is tough. I tend to solve this by only logic sooo. This might take a while

OpenStudy (anonymous):

a = 14 b = 35

OpenStudy (anonymous):

14 * 5 = 70 and 35 * 2 = 70.

OpenStudy (anonymous):

35-14 = 21, so it follows all guidelines.

OpenStudy (pawanyadav):

As per question | a-b|=21. (as difference is given a and b are not we not know who is greater) Again 5a-2b=18 Solve the two equations You will get the answer.